Trying to help my granddaughter solve a math problem. She's just now getting into geometry. Trying to solve the equation for an area of 130 square inches by 46 perimeters. What are the length and width of the triangle.
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

The rectangle is 10 wide by 13 long:
Perimeter = 2L+2W = 2(10)+2(13)= 20+26 = 46 Correct
Area = L*W = 10*13 = 130 Correct
```````````````````````````````````````````
You are missing information. I didn't need it because I knew the answer as soon as I saw the numbers, but I think what you are missing is something like "length is width plus 3" or something along those lines. Anyway, hopefully your granddaughter can use this answer (my granddaughter is 4, so I have a few years to go before I help her with math).
John

Answer by ikleyn(52781)

.
I think that yours triangle is actually a rectangle.

Then all information is in place, nothing missed, and the answer is:

The length is 13 and the width is 10.

The problem is very standard and is solved by reduction to a quadratic equation.

But first confirm that your triangle is a rectangle.

In this site we can help to your granddaughter and to you under one certain condition:

YOU WILL TRANSLATE THE PROBLEMS CORRECTLY.